PP.63 Jupiter, a large candy company, is having great success with its "Swan" family of candy bars. Due to a number of factors they like to plan their production at least nine months into the future. The table below contains their demand projections (in tons) for April through December:

Supply/Demand Info Beginning Apr May Jun Jul Aug Sep Oct Nov Dec
Predicted sales 3,000 2,200 1,900 2,000 2,400 3,200 4,800 2,700 3,600
Regular production
Overtime production
Subcontract production
Ending inventory 600
Hired employees
Fired employees
Total employees 80

Cost variables are as follows:

Cost Variables
Labor cost/hour $16
Overtime cost/ton $66
Subcontracting cost/ton $75
Holding cost/ton/month $10
Hiring cost/employee $1,000
Firing cost/employee $2,500

Here is some additional relevant (capacity) information:

Capacity Information
Total labor hours/ton 4
Regular production tons/employee/month 25
Max regular production/tons/month 3,000
Max overtime production/tons/month 1,500
Max subcontractor production/tons/month 2,000

Given the above information (and don't overlook beginning number of employees and inventory levels in the first table), create a level production plan with only the use of regular production and no inventory left over at the end of the nine-month period.

What is the regular production cost (over the nine months from April through December) for a level production plan? (Display your answer to the nearest whole number.)

What is the total overtime production cost for this production plan? (Display your answer to the nearest whole number.)

What is the total subcontract cost for this production plan? (Display your answer to the nearest whole number.)

What is the total holding cost for this production plan? (Display your answer to the nearest whole number.)

What is the total hire cost for this production plan? (Display your answer to the nearest whole number.)

What is the total fire cost for this production plan? (Display your answer to the nearest whole number.)

What is the total cost (sum of all costs) for this production plan? (Display your answer to the nearest whole number.)
Answer #1

Total predicted sales = 25800

Net production required = Total predicted sales - Beginning inventory = 25800-600 = 25200

Level production required / month = 25200/ 9 = 2800 (This is possible as max regular production is allowed 3000 tons)

Total employees required for 2800 tons/ month = 2800 /25 = 112

Please find below the production plan.

Ending inventory of the month = Regular production of the month + Overtime production of the month + Subcontract production of month + ending inventory of previous month - Predicted sales

Supply/Demand Info Beginning Apr May Jun Jul Aug Sep Oct Nov Dec Total
Predicted sales 3,000 2,200 1,900 2,000 2,400 3,200 4,800 2,700 3,600 25,800
Regular production 2800 2800 2800 2800 2800 2800 2800 2800 2800 25200
Overtime production 0 0 0 0 0 0 0 0 0 0
Subcontract production 0 0 0 0 0 0 0 0 0 0
Ending inventory 600 400 1,000 1,900 2,700 3,100 2,700 700 800 0 13300
Hired employees 32 0 0 0 0 0 0 0 0 32
Fired employees 0 0 0 0 0 0 0 0 0 0
Total employees 80 112 112 112 112 112 112 112 112 112
Regular production cost 179200 179200 179200 179200 179200 179200 179200 179200 179200 $1,612,800.00
Inventory cost = Ending inventory * 10 4000 10000 19000 27000 31000 27000 7000 8000 0 $   133,000.00
Hire cost $ 32,000.00 0 0 0 0 0 0 0 0 $     32,000.00
Total $1,777,800.00

a) Regular production cost =  Labor cost/hour * Total labor hours / ton * Tonnes produced* 9 months = 16*2800*4*9= $1612800

b) total overtime production cost = 0 (Since there is no overtime production)

c) Total subcontract cost = 0

d) Total holding cost= Holding cost /month * Total ending inventory =13300*10=$133000

e) Total hire cost = 32 * 1000 = $32000

f) Total fire cost = 0

g) total Cost = 1612800+133000+32000 = $1777800